# Physics-Informed Time-Integrated DeepONet: Temporal Tangent Space Operator Learning for High-Accuracy Inference

### v2

@@ 4 Discussion
PITI-DeepONet offers rapid, high-accuracy inference with limited to no training data as demonstrated by the numerical examples. Extrapolation using this method proves to be more stable and outperforms AR and FR methods, even extending far beyond the domain that was sampled for training: for the heat equation example up to 10 times the training domain. While the learned temporal tangent space operator incurs a higher inference cost compared to FR, it surpasses these methods in accuracy. We further explored different architectures for learning the temporal tangent space of time-evolution PDEs, but the results are similar to worse, while the presented architecture is the most general, being applicable to PDEs with mixed and/or higher-order derivatives in time and space. For the latter, one might incorporate an additional consistency loss, although the optimization process may become more challenging. While not the focus of this study, the accuracy of our model depends on the reliability of AD, and it may be affected if AD produces unstable results.
@@
A crucial aspect of effectively employing PITI-DeepONet is ensuring that all potential states encountered during inference are adequately sampled. These states can be obtained through experiments, traditional methods such as numerical or analytical solutions, or alternative surrogate models, such as FR simulations. Incorporating expert domain knowledge can significantly enhance the efficiency and quality of the sampling process. An additional, yet unexplored, avenue is the use of multi-fidelity learning [howard2023multifidelity], where data of varying resolutions or qualities can be seamlessly included within this framework. Moreover, it has been shown that a hybrid approach, which directly incorporates observed data, can greatly improve the learning process. However, this comes with the potential requirement of time derivative data for training, which may or may not be readily available, depending on the method used to acquire the data.
It is important to emphasize that no specific assumptions were made regarding the neural network architecture, allowing for the use of flexible, problem-specific designs. This can include architectures tailored to the number of variables, such as multi-input [jin2022mionet] and multi-output [Lu2022], as well as multi-network setups, i.e., a single network per field variable in coupled problems [Cai2021]. In this work, DeepONets with traditional feed-forward neural networks for the branch and trunk, as well as convolutional neural networks for the branch, were utilized; however, other (physics-informed) NO architectures are equally applicable. Furthermore, this framework can be readily scaled to higher-dimensional PDEs by adopting separable architectures [Mandl2025].
@@ 5 Conclusion
In summary, a practical architecture, PITI-DeepONet, is presented that achieves long-term accuracy with reasonable inference time, enabling near real-time predictions while incurring only a marginal increase in computational cost compared to traditional operator learning methods. The proposed approach leverages physics-informed temporal tangent space learning with restricted sampling, exhibiting strong, reliable extrapolation capabilities. Although the current study focused on three academic benchmark problems, this method has the potential to transform surrogate operator learning by bridging the gap between data-driven/physics-informed learning and classical numerical solvers. As a natural next step, the approach will be extended to tackle more challenging and potentially chaotic/stiff PDE systems.

### v3

@@ 3.5.1 Random Seed
3.5.2 Available Data and Initial Conditions
Figure 20: Data ablation study for Heat, Burgers’, Allen–Cahn Equation, and Kuramoto-Sivashinsky equation. Models are retrained with fixed hyperparameters while varying the fraction of available training data from $0.05$ to $1.0$ (step $0.05$). Solid lines show the mean relative $\mathcal{L}_{2}$ error over a fixed test set, and shaded regions indicate min/max across test examples. Inference uses explicit Euler time integration.
We retrain the models using the same reported hyperparameters and training budget as in the main experiments while varying the available training data fraction from $0.05$ to $1.0$ in increments of $0.05$ to quantify how performance scales with the number of training samples. All runs are evaluated on the same fixed test set and use explicit Euler time integration during inference. Figure 20 reports the mean relative $\mathcal{L}_{2}$ error across test examples together with the minimum and maximum error observed over the test set.
The results indicate that the sensitivity to additional training data is strongly benchmark-dependent. For the 1D heat equation, the mean error remains largely unchanged over the full range of data fractions, which is consistent with the diffusion-dominated character of this example and implies that the benchmark is saturated already at small fractions. While the mean performance changes only marginally, the min/max envelope becomes slightly tighter, indicating improved worst-case behavior with additional data but also decreased best-case performance. In contrast, for the 1D Burgers’ equation, the mean error decreases more noticeably with increasing data and essentially reaches the final performance already at approximately $60\%$ of the training data, after which further gains are small. Lastly, the error for the higher-dimensional Allen–Cahn example shows signs of convergence toward bigger fractions and drops more gradually over the whole range, reflecting the increased complexity of the dynamics. This response is more pronounced in KS equation, which even indicates that further improvement seems achievable with more data than used within this work.
For hybrid training configurations, data-driven components are scaled consistently with the available data fraction. Overall, the ablation confirms that the proposed method remains robust even when the training data is significantly reduced.
3.5.3 Inference Time Step
Figure 21: Inference time step analysis using explicit Euler. The models are used in inference with varying step size $\Delta t$. Solid lines show the mean relative $\mathcal{L}_{2}$ error across test examples, and shaded regions indicate min/max across the test set.
Since the proposed approach advances the solution by time integrating a learned temporal tangent operator, we assess robustness with respect to the inference step size $\Delta t$. We note that the manuscript already includes a time-step study for inference in Fig. 10, where we used a smaller and larger $\Delta t$ for the Burgers’ equation, while all other inference results in the manuscript were obtained with $\Delta t=0.01$. Building on this, we sweep the inference step size $\Delta t$ over several orders of magnitude using explicit Euler, without retraining the models, and evaluate all runs on the same fixed test set. We report the minimum, mean, and maximum relative $\mathcal{L}_{2}$ error across test examples (cf. Fig. 21).
Across all four PDEs, we observe a plateau where decreasing $\Delta t$ does not lead to increased accuracy. This suggests that once $\Delta t$ is sufficiently small, the learned tangent operator is the primary source of error rather than temporal discretization. For coarser steps, the error increases due to loss of temporal resolution, with a sharper deterioration for the heat equation and a more gradual deterioration for Burgers’ equation, consistent with the need to resolve stronger nonlinear/advection-driven dynamics. For Allen-Cahn, the mean error slightly improves for larger $\Delta t$, which is in line with reduced accumulation of small systematic model biases when fewer Euler steps are taken, especially because smaller $\Delta t$ implies many more Euler steps, so small systematic errors in the learned tangent can accumulate over time. Nonetheless, for the Allen–Cahn example, the error is largely invariant with respect to the chosen time step $\Delta t$ over the tested range. Lastly, this behavior is also present for the KS equation, as the method would benefit strongly from a better learned derivative than using a smaller timestep.
We emphasize that this experiment is intended as a step-size robustness study for the learned dynamics, rather than a classical solver-order convergence test, because the learned operator introduces an error floor and repeated operator evaluations can accumulate model bias when $\Delta t$ becomes very small. Furthermore, the time step $\Delta t=0.01$ in the main part is a good choice for all PDEs, with the potential to use higher time steps for the heat and Allen-Cahn equations. Finally, the absolute error levels differ across PDEs, reflecting differences in the difficulty of learning the temporal tangent operator for the respective state distributions, rather than differences in the time-step size being the limiting factor.
@@ 4 Discussion
PITI-DeepONet offers rapid, high-accuracy inference with limited to no training data, as demonstrated by the numerical examples even for equations that are considered chaotic. Extrapolation using this method proves to be more stable and outperforms AR and FR methods, even extending far beyond the domain that was sampled for training: for the heat equation example up to 10 times the training domain. While the learned temporal tangent space operator incurs a higher inference cost compared to FR, it surpasses these methods in accuracy. We further explored different architectures for learning the temporal tangent space of time-evolution PDEs, but the results are similar or worse, while the presented architecture is the most general, applying to PDEs with mixed and/or higher-order derivatives in time and space. For the latter, one might incorporate an additional consistency loss, although the optimization process may become more challenging. While not the focus of this study, the accuracy of our model depends on the reliability of AD, and it may be affected if AD produces unstable results.
@@
A crucial aspect of effectively employing PITI-DeepONet is ensuring that all potential states encountered during inference are adequately sampled. These states can be obtained through experiments, traditional methods such as numerical or analytical solutions, or alternative surrogate models, such as FR simulations. Incorporating expert domain knowledge can significantly enhance the efficiency and quality of the sampling process. An additional, yet unexplored, avenue is the use of multi-fidelity learning Howard et al. [2023], where data of varying resolutions or qualities can be seamlessly included within this framework. Moreover, it has been shown that a hybrid approach, which directly incorporates observed data, can greatly improve the learning process. However, this comes with the potential requirement of time derivative data for training, which may or may not be readily available, depending on the method used to acquire the data.
It is important to emphasize that no specific assumptions were made regarding the neural network architecture, allowing for the use of flexible, problem-specific designs. This can include architectures tailored to the number of variables, such as multi-input Jin et al. [2022] and multi-output Lu et al. [2022], as well as multi-network setups, i.e., a single network per field variable in coupled problems Cai et al. [2021]. In this work, DeepONets with traditional feed-forward neural networks for the branch and trunk, as well as convolutional neural networks for the branch, were utilized; however, other (physics-informed) NO architectures are equally applicable. Furthermore, this framework can be readily scaled to higher-dimensional PDEs by adopting separable architectures Mandl et al. [2025].
@@ 5 Conclusion
In summary, a practical architecture, PITI-DeepONet, is presented that achieves long-term accuracy with reasonable inference time, enabling near real-time predictions while incurring only a marginal increase in computational cost compared to traditional operator learning methods. The proposed approach leverages physics-informed temporal tangent space learning with restricted sampling, exhibiting strong, reliable extrapolation capabilities. Although the current study focused on four academic benchmark problems, this method has the potential to transform surrogate operator learning by bridging the gap between data-driven/physics-informed learning and classical numerical solvers. As a natural next step, the approach will be extended to tackle more challenging and potentially chaotic/stiff PDE systems.